OpenStudy (anonymous):

Two lines, A and B, are represented by the following equations: Line A: 4x + 4y = 16 Line B: x + y = 4 Which statement is true about the solution to the set of equations? It is (1, 2). There are infinitely many solutions. It is (1, 5). There is no solution.

OpenStudy (campbell_st):

what happens when you divide every term in line A by 4...?

OpenStudy (carolinar7):

Infinite

OpenStudy (anonymous):

Consider the following set of equations: Equation R: −3y = −3x – 9 Equation S: y = x + 3 Which of the following best describes the solution to the given set of equations? No solution One solution Infinite solutions Two solutions

OpenStudy (carolinar7):

because they are the same

OpenStudy (carolinar7):

Infinite

OpenStudy (carolinar7):

Because they are the same

OpenStudy (anonymous):

Consider the following equations: −x − y = 1 y = x + 3 If the two equations are graphed, at what point do the lines representing the two equations intersect? (−1, 2) (−2, 1) (1, −2) (2, −1)

OpenStudy (carolinar7):

Just plug in the values and see which one works

OpenStudy (carolinar7):

SO (-1,2) ,(-2,1)

OpenStudy (anonymous):

I don't know I don't get the concept

OpenStudy (carolinar7):

I ts just substituting in values

OpenStudy (anonymous):

1,-2

OpenStudy (carolinar7):

When you substitue in that value do both side equal the same.

OpenStudy (anonymous):

-1,2

OpenStudy (carolinar7):

try to plu that in y=x+3

OpenStudy (carolinar7):

ya -1,2 works

OpenStudy (anonymous):

Consider the following set of equations: Equation A: y = 2x + 4 Equation B: y = 3x + 1 Which of the following is a step that can be used to find the solution to the set of equations? 2x + 4 = 3x + 1 2x = 3x + 4 2x + 4 = 3x 2x + 1 = 3x + 4

OpenStudy (anonymous):

I got 2x+4=3x+1

OpenStudy (carolinar7):

ya

OpenStudy (anonymous):

is that right

OpenStudy (carolinar7):

ya

OpenStudy (anonymous):

Consider the following system of equations: y = −x + 2 y = 3x + 1 Which description best describes the solution to the system of equations? Line y = −x + 2 intersects line y = 3x + 1. Lines y = −x + 2 and y = 3x + 1 intersect the x-axis. Lines y = −x + 2 and y = 3x + 1 intersect the y-axis. Line y = −x + 2 intersects the origin.

OpenStudy (anonymous):

Lines y = −x + 2 and y = 3x + 1 intersect the y-axis this is what I got

OpenStudy (carolinar7):

I got none of those

OpenStudy (anonymous):

what did you get

OpenStudy (carolinar7):

I got at (.25,1.75)

OpenStudy (anonymous):

Consider the following pair of equations: y = x + 4 y = −2x − 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (carolinar7):

Doing Module 6 test

OpenStudy (anonymous):

no module 8

OpenStudy (carolinar7):

Oh Algebre I

OpenStudy (carolinar7):

or II

OpenStudy (anonymous):

I think 1

OpenStudy (carolinar7):

How old

OpenStudy (carolinar7):

are you

OpenStudy (anonymous):

16

OpenStudy (carolinar7):

oh

OpenStudy (carolinar7):

either way Anser is (-2,2)

OpenStudy (anonymous):

I need to show my work how to get tnswer

OpenStudy (carolinar7):

OK

OpenStudy (carolinar7):

1.Substitute -2x-2=x+4 2. Solve for x. -2x-2=x+4 +2x +2x -2=3x+4 -4 -4 -6= 3x /3 /3 -2=x 3. Now find the y value by substituting in the x value

OpenStudy (anonymous):

cool pic

OpenStudy (carolinar7):

2=(-2)+4 so (-2,2)

OpenStudy (anonymous):

Not here to help you with math, just saying yo

OpenStudy (anonymous):

I have one more test

OpenStudy (carolinar7):

Hey Stop trolling and wasting my time i=on purpose